This document shows an example of the use of the subeqnarray environment. Here is one:

x = a × b  (1a)
= z + t  (1b)
= z + t  (1c)

The first equation is number 1a, the last is 1c. The equation as a whole can be referred to as equation 1.

To show that equation numbers behave normally, here’s an eqnarray environment.

x = a × b  (2)
= z + t  (3)
= z + t  (4)

These are equations 2, 3 and 4.